OpenStudy (campayne):

Use synthetic substitution to evaluate P(x) = x^3 + 3x^2 - 6 for x = -1. P(-1) = ____ Enter a number only.

OpenStudy (bonnieisflash1.0):

need some help

OpenStudy (campayne):

quite a bit

OpenStudy (bonnieisflash1.0):

or nah

OpenStudy (campayne):

yeah im kinda dying here

OpenStudy (bonnieisflash1.0):

what you mean?

OpenStudy (campayne):

im dying i need help

OpenStudy (bonnieisflash1.0):

okay

OpenStudy (bonnieisflash1.0):

hey @ahsan are you new here

OpenStudy (bonnieisflash1.0):

here If one root is -3, the remainder will be zero when the polynomial is divided by (x+3). Using synthetic division, we can evaluate the remainder easily. .. 2, -3, p, 30 ... first row of synthetic division is the coefficients of the polynomial .. - , -6, 27, (-3p-81) ... result of multiplying the bottom row by -3 .. 2, -9, (p+27), (-3p-51) ... bottom row is the sum of the first two rows. In order for the remainder to be zero, the value of (-3p-51) must be zero. .. -3p - 51 = 0 .. -3p = 51 .. p = 51/-3 = -17 The coefficients of the polynomial resulting from division by (x+2) are {2, -9, -17+27}, so the original polynomial factors as .. (x+3)(2x^2-9x+10) = 0 .. (x+3)(2x-5)(x-2) = 0 ... factoring the quadratic The value of p is -17, and the other roots are 2.5 and 2.

OpenStudy (bonnieisflash1.0):

and bye

OpenStudy (campayne):

wut

OpenStudy (bonnieisflash1.0):

huh?

OpenStudy (campayne):

i dont get it

OpenStudy (bonnieisflash1.0):

okay